# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

### v1

@@ 5.3 GTTA Classification Results
Results on CIFAR-10C, CIFAR-100C, and Imagenet-C. This subsection presents the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C. For each task, the corruption severity of the inputs is gradually increased from levels 1 to 5, as discussed in the implementation details in Sec. 5.1. Table 4 reports the mean classification error for GTTA across all severities. We observe that, for all the datasets, our approach outperforms or performs comparably well across all the baselines [37, 44, 45, 3, 33]. With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA [45], which also demonstrates competitive improvements in the mean classification error.
@@ 5.4 Ablation Studies
Figure 2: Illustrations of the learning rate importances of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10] during CTTA, for different convolutional blocks/stages/layers. [Top row] - Variations across respective datasets for domain glass_blur. [Bottom row] - Variations across respective datasets for domain snow.
Analysis of the learning rate importance. The prime objective of our work is to adaptively alter the learning rates of parameters of selected layers based on their sensitivity. In Fig. 2, we illustrate how the learning rate (LR) importance (Eq. 11) of WideResNet-28 [48], ResNeXt-29 [46], and ResNet-50 [10], averaged for a task, change during adaptation on their respective datasets. For brevity, we show two domains - glass_blur and snow. As we can see, the importance varies across domains and models. Across all the models, we notice that generally, the deeper layers have smaller associated LR importances, which means that the initial layers are adapted more to the target domain. This result completely aligns with the findings of surgical fine-tuning [21]. In addition to that, we also notice more sparsity in the deeper layers. This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
Nullifying prediction uncertainty. In LAW [33], the authors make use of the pseudo-labels to automatically decide on the importance of layers for adaptation. In this, we study the effectiveness of such a design for our method i.e. we investigate if layer importance computed using pseudo-labels is helpful. We compute the gradients from the log-likelihood (Eq. 2) and estimate the sensitivity of the parameters of the chosen layers. Table 5.4 reports the mean classification errors for all the datasets. We observe that the gradients computed using the pseudo-labels are an unreliable indicator of prediction uncertainty. The scoring function (Eq. 6) computed using these gradients carries insufficient information. We achieve improvements of 2.44%, 3.69%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
@@
Table 5: Prediction uncertainty using pseudo-labels vs our method. Dataset Pseudo-labels Ours CIFAR-10C 17.98 15.54 CIFAR-100C 33.88 30.19 ImageNet-C 68.62 60.12
Table 6: Ablation results on benchmark datasets with varying $\lambda$. $\lambda$ 0 0.01 0.05 0.1 0.5 1.0 CIFAR-10C 16.84 15.54 16.45 17.03 17.65 17.93 CIFAR-100C 30.44 30.19 30.51 30.57 30.63 30.7 ImageNet-C 63.98 60.12 60.38 60.4 60.4 60.42
Contribution of smoothing factor $\alpha$. An important component of our method is the computation of the moving average of sensitivities across data batches, smoothed by a parameter $\alpha$ (Eq. 9). Fig. 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the "domain-level" sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e. when the current sensitivity of parameters of selected layers is not considered, the mean classification error for the respective datasets drastically spikes up. This suggests that the effect of the current sensitivity of selected layers is indeed required. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.67% over LAW [33]. Similarly, for CIFAR-100C and ImageNet-C, we see improvements of 1.56% and 1% over LAW [33], at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much and are stable. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{n}^{t}$) and current sensitivity ($S_{n}^{t}$) is critical, as seen in the figure.
@@
Robustness to the regularisation coefficient $\lambda$. As an important ablation step, we understand the effect of the parameter $\lambda$ on the final optimization objective, as in Eq. 16. We set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}. We study the contributions of the consistency loss to entropy minimization. Table 5.4 reports the average classification errors on CIFAR-10C, CIFAR-100C, and ImageNet-C, with varying $\lambda$. For $\lambda$ = 0.01, we achieve the best results across all the datasets. In addition, we observe that our method is very robust to $\lambda$ i.e. with subtle variations to differing $\lambda$, across the datasets. In particular, for CIFAR-10C, with $\lambda$ = 0 i.e. with no consistency loss, we achieve improvements of 0.37% and 3.86% over LAW [33] and TENT Continual [44]. For CIFAR-100C, with no consistency loss ($\lambda$ = 0), we notice improvements of 1.31%, 2.06%, 4.96%, and 30.46% over LAW [33], CoTTA [45], BN Stats Adapt [37] and TENT Continual [44]. As for ImageNet-C, the consistent best setting of $\lambda$ = 0.01 is better than LAW [33] by 1% and by 5.38% over Adacontrast [3]. However, with $\lambda$ = 0 i.e. no consistency loss, we see a spike in the error rate by about 3.86% over the best setting ($\lambda$ = 0.01).
@@ 6 Conclusion
In this paper, we propose a novel approach called PALM, to continual test-time adaptation. Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e. by backpropagating the KL divergence between a uniform distribution and the softmax outputs. This effectively eliminates the need for pseudo-labels. For the parameters of these layers, selected based on the norm of the gradients, the learning rates are adjusted based on their sensitivity to domain shift. Extensive experiments are conducted on benchmark datasets to demonstrate the efficacy of our method against different baselines. The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.

### v2

@@ 5.3 GTTA Classification Results
Results on CIFAR-10C, CIFAR-100C, and Imagenet-C. Here, we present the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C. For each task, the corruption severity of the inputs is gradually changed as discussed earlier. Table 4 reports the mean classification error for GTTA across all severities. We observe that, for all the datasets, our approach outperforms or performs comparably well across all the reported baselines. With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA, which also demonstrates competitive improvements in the mean classification error.
@@ 5.4 Ablation Studies
Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA. [Top row] - Variations for the domain “glass_blur" across datasets. [Bottom row] - Variations for the domain “snow".
Dataset Method
The average % of
adapted parameters/task
(source parameters)
CIFAR-10C Ours 4.34 (of 36.4M)
CoTTA 100 (of 36.4M)
CIFAR-100C Ours 6.65 (of 6.8M)
CoTTA 100 (of 6.8M)
ImageNet-C Ours 17.21 (of 23.5M)
CoTTA 100 (of 23.5M)
Table 5: Average % of adapted parameters/task.
Computational efficiency and Catastrophic forgetting: In our proposed method, we first select a subset of layers from the gradient space (Eq. 2) and adapt them while strictly freezing the unselected layers. In Table 5, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets. Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source model parameters. This showcases significant computational advantages over CoTTA, which updates 100% of the source parameters. Additionally, in LAW, the LRs of all the parameters are adjusted. It is worth mentioning that our method alleviates catastrophic forgetting (CF) by only adapting a small subset of model parameters ensuring that much of the prior source knowledge is preserved and CF is largely minimized.
Dataset Ours LARS Prodigy Adafactor
CIFAR-10C 15.5 18.9 20.4 73.8
CIFAR-100C 30.1 34.1 35.5 97.7
ImageNet-C 60.1 62.0 68.6 99.2
Table 6: CTTA mean errors (%) using adaptive LR optimizers vs. ours.
Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS [41], Prodigy [19], and Adafactor [32], for CTTA. We report the results in Table 6. The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
Analysis of LR importance $i_{j,n}^{t}$: The prime objective of our work is to adaptively alter the LRs of parameters of selected layers based on their sensitivity. In Figure 2, we illustrate how the LR importance $i_{j,n}^{t}$ (Eq. 7) of WideResNet-28, ResNeXt-29, and ResNet-50, averaged for a task, change during adaptation on their respective datasets. For brevity, we show two domains - glass_blur and snow. We observe that the importance varies across domains and models. Across all the models, we notice that generally, the deeper layers have smaller associated $i_{j,n}^{t}$, which means that the initial layers are adapted more to the target domain. This result completely aligns with the findings of surgical fine-tuning [15]. In addition to that, we also notice more sparsity in the deeper layers. This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
Layer importance via pseudo-labels: In LAW, the authors use the pseudo-labels to decide on the importance of layers for adaptation. In this, we study the effectiveness of such a design for our method i.e., we investigate if layer importance computed using pseudo-labels is helpful. We compute the gradients from the log-likelihood as $\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$ and estimate the sensitivity of the parameters of the chosen layers (Eq. 4-Eq. 6), and alter the learning rates (Eq. 7). Table 7 reports the mean classification errors for all the datasets. We observe that the gradients computed using the pseudo-labels are an unreliable indicator of prediction uncertainty. The scoring function $\mathcal{Z}_{\theta_{j,n}^{t}}=\|{{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}\|_{1}$ computed using these gradients carries insufficient information. We achieve improvements of 2.4%, 3.7%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
@@
Dataset Pseudo-labels Ours
CIFAR-10C 17.9 15.5
CIFAR-100C 33.8 30.1
ImageNet-C 68.6 60.1
Table 7: Layer importance using pseudo-labels vs ours.
$\lambda$ 0 0.01 0.05 0.1 0.5 1.0
CIFAR-10C 16.8 15.5 16.4 17.0 17.6 17.9
CIFAR-100C 30.4 30.1 30.5 30.5 30.6 30.7
ImageNet-C 63.9 60.1 60.3 60.4 60.4 60.4
Table 8: Ablation results with varying $\lambda$.
Contribution of smoothing factor $\alpha$: A key component of our method is computing the moving average of sensitivities across data batches, smoothed by parameter $\alpha$ (Eq. 5). Fig. 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the “domain-level" sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes. This suggests that the effect of the current sensitivity of selected layers is indeed required. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW. For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
@@
Robustness to regularization $\lambda$: Here, we understand the effect of the parameter $\lambda$ in Eq. 11. We set $\lambda$in {0, 0.01, 0.05, 0.1, 0.5, 1.0}. Table 8 reports the average classification errors on the benchmark datasets. For $\lambda$ = 0.01, we achieve the best results. In addition, we observe that our method is very robust to $\lambda$ i.e., with subtle variations to differing $\lambda$.
@@ 6 Conclusion
In this paper, we propose a novel approach called PALM, for continual test-time adaptation. Our approach automatically decides on the importance of a layer by leveraging information from the gradient space i.e., by backpropagating the KL divergence between a uniform distribution and the softmax outputs. This effectively eliminates the need for pseudo-labels. For the parameters of these layers, selected based on the norm of the gradients, the learning rates are adjusted based on their sensitivity to domain shift. Extensive experiments are conducted on benchmark datasets to demonstrate the efficacy of our method against different baselines. The results reveal that PALM achieves superior performance and pushes the state-of-the-art on several standard baselines.